1. Tungsten carbide (WC), used in making tools and jewelries, is produced by the reaction of tungsten hexachloride (WCl6), H2 and CH4 at 670℃.

WCl6 + H2 + CH4 WC + 6HCl


If 23.5 g WCl6 reacts with excess H2 and CH4, and 10.8 g WC are recovered, what is the percent yield?



1
Expert's answer
2021-11-17T18:29:01-0500

Taking W=183.84 Cl=35.453 C=12.0107


WCl6= 183.84+ (35.453×6)= 396.558g


WC= 183.84+ 12.0107

= 195.8507g


396.558 of WCl6 yields 195.8507 of WCl6


23.5g of WCl6 in the same proportion should yield


"=\frac{23.5}{396.558}×195.8507=11.6061g"


of WCl6


Actual yield given = 10.8g


"\%" yield "=\frac{Actual \>yield}{Theoretical \> value}×100\%"



"=\frac{10.8}{11.6061}×100\%=93.06\%"